OpenStudy (mergirlblue):

what is s_10 for 1+3+9+27 + ....?

OpenStudy (michele_laino):

hint: here we have a geometri series, whose constant \(r\) is: \[\Large r = \frac{{27}}{9} = \frac{9}{3} = \frac{3}{1} = 3\]

OpenStudy (michele_laino):

geometric*

OpenStudy (michele_laino):

now, the requested sum \(S_{10}\), is: \[\Large {S_{10}} = {a_1} \cdot \frac{{{3^{10}} - 1}}{{3 - 1}}\] where \(a_1=1\) is the first term of the series

OpenStudy (michele_laino):

then we can write: \[\Large {S_{10}} = {a_1} \cdot \frac{{{3^{10}} - 1}}{{3 - 1}} = 1 \cdot \frac{{{3^{10}} - 1}}{{3 - 1}} = ...?\]

OpenStudy (mergirlblue):

29,524 ?

OpenStudy (mergirlblue):

Thank you

OpenStudy (michele_laino):

that's right! :)